OpenStudy (anonymous):

what is demosthenes had a bag of pebbles when the sun came up.he put 17 pebbles in his mouth. then there were 46 pebbles in the bag how many pebbles were in the bag when thre sun came up?write a subtraction pattern and solve the problem.

OpenStudy (anonymous):

Let's call the amount of pebbles in the bag when the sun came up x. He put 17 in his mouth and then there were still 46 left. so when we subtract 17 from x we get 46. Can you write down the equation now?

OpenStudy (anonymous):

17-x=46 yes

OpenStudy (anonymous):

Are you sure about that? We're supposed to subtract 17 from x.

OpenStudy (anonymous):

can you teach me plz? again

OpenStudy (anonymous):

Is it x-17=46 or 17-x=46? what do you think?

OpenStudy (anonymous):

oh okay thanks i will help other people now

OpenStudy (anonymous):

i will help people now

OpenStudy (anonymous):

thanks